OpenStudy (mstv1112):

4 questions. Fan and Medal :)

OpenStudy (mstv1112):

9. State whether the equation is true, false, or an open sentence. Show your work. 3(4 + –5) = 3(–6 – 2)

OpenStudy (mstv1112):

8. Name the coefficients in the expression 4x + 9 – y

OpenStudy (mstv1112):

7. Marcus buys groceries for no more than $250.00 each week. He shops 4 times each month. How much should he budget for groceries for the month?

OpenStudy (mstv1112):

\[\frac{ y }{ 4 }\le-16\] Solve for y please show steps

OpenStudy (anonymous):

9. 3(-1)=3(-6-2) -3=3(-8) -3=-24 21=0. It can not be that.

OpenStudy (anonymous):

8. coefficients:4,9,1

OpenStudy (anonymous):

7. $1000 6.\[y \le-64\]

OpenStudy (mstv1112):

How did you get the last one?

OpenStudy (mstv1112):

@charliealbert

OpenStudy (anonymous):

So in order to isolate y. We have to multiply both sides by for. So we are going to obtain: \[4(\frac{ y }{ 4 } \le16)=\frac{ 4 }{ 4 }y \le16*4=y \le64\]